OpenStudy (anonymous):

In a given amount of time, Mary drove twice as far as Lynn. Altogether they drove 60 miles. How far did Mary drive? A. 20 miles B. 30 miles C. 40 miles D. 45 miles

OpenStudy (btaylor):

Using M for the distance Mary drove and L for the distance Lynn drove, how would you write the "Mary drove twice as far as Lynn" as an equation? What about the second part, "together they drove 60 miles"? Once you can convert these to equations, you should be able to solve the system by substitution.

OpenStudy (anonymous):

B ?

OpenStudy (btaylor):

@DUDE..IM..A..DUCK please try to solve the problem and not just guess answers.

OpenStudy (nirmalnema):

let the car driven by lyn=S there by mary = 2s s+2s=60 3s=60 s=20 so car driven by masy = 2*20=40

OpenStudy (anonymous):

I figured 30 and 2 would equal the equation like 1.21 gigawats , and a minimum of 80 mph to go back in time?

OpenStudy (anonymous):

I'd let you guys in a secret , but then that would ruin my fun